# Causal Judge Evaluation: Calibrated Surrogate Metrics for LLM Systems

### v2

@@ 5 Experiments
We evaluate CJE on a large-scale benchmark. We initially expected doubly robust methods to dominate by combining logged data with fresh draws. Instead, we found two surprises: (i) OPE fails catastrophically even after weight stabilization boosts ESS above 90%, and (ii) DR provides no advantage over Direct under low overlap, effectively reducing to the outcome model alone. These findings motivated the CLE diagnostic, which explains why high ESS is insufficient.
@@
    Does calibration work? AutoCal-R should reduce RMSE, improve ranking accuracy, and with OUA dramatically improve CI coverage.
@@ 5.1 Experimental design
Oracle $Y$: gpt-5-2025-08-07 (OpenAI, 2025) quality scores (0–100, normalized to $[0,1]$) collected for every response. Judge $S$: gpt-4.1-nano-2025-04-14 scores on all responses (temperature 0); approximately $16\times$ cheaper than oracle. [^1] Pricing as of November 2025 via OpenAI API. Note: The oracle model operates at temperature 1.0 (temperature 0 was not available for this model at experiment time), introducing stochasticity in oracle labels; we did not repeat-score responses, so coverage results may be conservative. All estimators except naive-direct calibrate judge scores using AutoCal-R. Base variants use monotone-only calibration (Z=SZ{=}S); +cov variants use two-stage calibration with response length as a covariate ($Z{=}g(S,\text{length})$). For DR methods, rewards $R=\hat{f}_{\mathrm{all}}(Z)$ use the full calibrator while the outcome model $\hat{q}=\hat{f}^{(-k)}(Z)$ uses cross-fitted (leave-one-fold-out) predictions; the difference $R-\hat{q}$ captures residual information for the IPS correction.
@@ 5.2 Results
Table 1: Accuracy Metrics
Estimator $\mathrm{RMSE}^{\mathrm{d}}$ $\downarrow$ Pairwise % $\uparrow$ Top-1 % $\uparrow$ $\tau$ $\uparrow$ Runtime (s) $\downarrow$
naive-direct 0.0828 90.9 79.6 0.817 0.4
direct 0.0223 91.8 84.1 0.837 0.7
direct+cov 0.0256 94.4 89.5 0.888 1.7
SNIPS 0.1596 38.3 8.7 -0.235 0.3
SNIPS+cov 0.1842 37.4 6.4 -0.252 0.3
calibrated-ips 0.0246 47.1 19.0 -0.059 0.3
calibrated-ips+cov 0.0262 46.5 20.5 -0.070 0.4
dr-cpo 0.0459 78.3 46.3 0.566 3.4
dr-cpo+cov 0.0570 79.1 50.1 0.581 10.4
calibrated-dr-cpo 0.0224 90.9 81.0 0.818 3.4
calibrated-dr-cpo+cov 0.0260 94.1 87.7 0.882 10.2
stacked-dr 0.0224 91.9 83.1 0.838 9.3
stacked-dr+cov 0.0281 92.1 84.8 0.842 42.7
$\downarrow$: lower is better, $\uparrow$: higher is better. Bold: best, underlined: second-best. Metrics averaged across all regimes. See Table 3 for uncertainty quantification.
@@
Yes, for both accuracy and inference. Table 1 shows calibration improves accuracy across method families. For Direct: uncalibrated naive-direct achieves 90.9% pairwise vs. calibrated direct+cov at 94.4%. For DR: raw-weight dr-cpo achieves 78.3% pairwise ($\tau{=}0.57$) vs. weight-stabilized calibrated-dr-cpo at 90.9% ($\tau{=}0.82$), a 12.6 percentage point gain (both use AutoCal-R; the gain is from SIMCal-W). For inference: uncalibrated CIs (naive-direct) achieve 0% coverage: nominal 95% CIs capture the true value in 0/50 seeds. Bootstrap inference with bias-corrected estimation ($\hat{\theta}_{\mathrm{aug}}$) dramatically improves coverage: Direct achieves $\sim$}95% across oracle fractions (Table 3), as does stacked-DR. The OUA share of total variance ranges from 0% (100% oracle) to over 50% (5% oracle), explaining why ignoring calibration uncertainty produces catastrophically narrow CIs (Figure 6).
@@
We expected DR to outperform Direct by combining logged data with fresh draws. Instead, direct+cov slightly outperforms calibrated-dr-cpo+cov (94.4% vs. 94.1%). The explanation: under low overlap, DR’s IPS component contributes noise rather than information. DR estimators are consistent if either the propensity model or the outcome model is correct. Given IPS’s failure, DR effectively relies entirely on the outcome model $\hat{q}(Z)$, not the propensities. This is exactly the regime CLE identifies: TTC is low, so logged data contributes little; DR $\approx$ Direct ++ extra variance from unstable weights. Implication: When teacher forcing is unreliable, prefer Direct over DR; it achieves the same accuracy without requiring logprobs.
@@
Effect of AutoCal-R (Reward Calibration)
@@
Direct + AutoCal-R 0.023 95% –72%
Effect of SIMCal-W (Weight Stabilization)
@@
DR-CPO, raw weights 0.046 99%
DR-CPO + SIMCal 0.023 100% –50%
DR-CPO + SIMCal + cov 0.026 99% –43%
@@
Direct + AutoCal-R + cov 0.026 95% –69%
Key findings: AutoCal-R reduces RMSE by 72% and restores coverage (0%→95% with bootstrap). SIMCal-W reduces IPS RMSE by 84%. Both contribute additively.
Table 3: Coverage by UQ Method (target: 95%). Bootstrap with calibrator refit achieves near-nominal coverage; cluster-robust and jackknife-only OUA undercover.
@@
Regime Cluster-Robust OUA Jackknife B=500B{=}500 B=2000B{=}2000
@@
Table 2 isolates the marginal contribution of each CJE component. AutoCal-R (reward calibration) reduces RMSE by 72% and is necessary to avoid catastrophic undercoverage: without it, 0% of CIs contain the true value; with bootstrap inference, coverage improves to $\sim$}95%. SIMCal-W (weight stabilization) reduces IPS RMSE by 84% while maintaining coverage, validating the $S$-monotone projection approach. Table 3 compares UQ methods for Direct mode: naive cluster-robust SEs achieve only 15–52% coverage; OUA jackknife improves to 70–89%; bootstrap with $\hat{\theta}_{\mathrm{aug}}$ (AIPW-style residual augmentation) achieves $\sim$}95% by correcting both variance and bias from calibration uncertainty.
Table 4: Weight Diagnostics: SIMCal Stabilization. Raw SNIPS weights have catastrophic ESS ($1\%$) and heavy tails ($\alpha<2$); SIMCal projection recovers ESS $80\%$ with light tails.
@@
Policy SNIPS SIMCal SNIPS SIMCal TTC CLE
@@
Figure 5: CJE output: policy value estimates at n=1000n{=}1000, 25% oracle. Red diamonds show oracle ground truth; blue circles show CJE estimates with 95% CIs. CIs capture the true value for policies satisfying transportability; unhelpful (which violates transportability) shows slight miscoverage, a failure mode flagged by the transportability test (Figure 7). This is representative of what a practitioner would see when applying CJE to their own evaluation data.
Figure 6: Uncertainty decomposition at n=1000n{=}1000 for direct+cov. Yellow: base sampling uncertainty (constant $\sim$0.014). Orange: oracle uncertainty (dominates at 5% oracle fraction, vanishes at 100%). At 5% oracle fraction, OUA contributes 55% of total variance; ignoring it produces invalid CIs.
@@
Figure 7 applies the policy-wise mean transport test (3) to validate whether the base-trained calibration can be reused across target policies. We test $H_{0,\pi^{\prime}}:\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$ for each policy with Bonferroni correction ($\alpha{=}0.0125$). Clone, Parallel Universe, and Premium all pass: their residual distributions are centered at zero, implying that Direct + calibrated surrogate remains mean-unbiased for these policies. Unhelpful fails dramatically (mean residual 0.31-0.31, $p<0.001$). By 3, this implies $V^{\mathrm{oracle}}_{\pi^{\prime}}-V^{\mathrm{sur}}_{\pi^{\prime}}\approx-0.31$: the surrogate systematically overestimates oracle quality for adversarial responses. This flags a failure of mean transportability: absolute value estimates for this policy are biased, triggering the REFUSE-LEVEL gate (App. D.6). Practical implication: This test decides whether a single calibration can be safely reused across policies; if a policy fails, either recalibrate with policy-specific oracle data or fall back to oracle-only evaluation for that cell.
@@
Figure 8: Sample size planning for direct+cov. (A) RMSE by sample size and oracle fraction. (B) Standard errors including OUA. (C) MDE contours showing the smallest effect detectable at 80% power. Dashed lines show common detection thresholds; cells with MDE <0.02<0.02 enable detecting 2-point quality differences. Example: To detect a 3% improvement with 80% power requires $n{\geq}1{,}000$ with ${\geq}10\%$ oracle fraction.
@@
Calibrated IPS Required + SIMCal-W stabilization
@@ 5.3 Practical takeaways
    Default to Direct + two-stage calibration. If you can generate fresh responses on a shared prompt set, use the Direct Method. It requires no overlap assumption and achieves the best ranking accuracy (Table 6).
@@
    Check TTC before using OPE. If TTC <0.7<0.7, logs-only IPS will fail. Use DR only when fresh draws are available to complement logged data.
@@
    Always use bootstrap inference with $\hat{\theta}_{\mathrm{aug}}$. Standard CIs severely under-cover (15–52%); OUA jackknife improves but still falls short (70–89%); bootstrap with bias correction achieves near-nominal coverage ($\sim$}95%).
@@
    Use OUA share to guide resource allocation. The fraction $\widehat{\mathrm{Var}}_{\mathrm{cal}}/\widehat{\mathrm{Var}}_{\mathrm{total}}$ identifies bottlenecks: if OUA share $50\%$, collect more oracle labels; if OUA share $20\%$, collect more evaluation prompts (Figure 6).
@@
  6. 6.
    Optimize the budget ratio. How many oracle labels do you need? The optimal ratio of labels ($m$) to scores ($n$) follows a square-root law based on costs and variances (Appendix F). Check the OUA share $\omega$ to equalize marginal utility: if $\frac{\omega}{1-\omega}>\frac{\mathrm{Spend}_{\mathrm{oracle}}}{\mathrm{Spend}_{\mathrm{surrogate}}}$, the experiment is under-investing in oracle labels.

### v3

@@ 5 Experiments
We evaluate CJE on a large-scale benchmark. We initially expected doubly robust methods to dominate by combining logged data with fresh draws. Instead, we found two surprises: (i) OPE fails catastrophically even after weight stabilization boosts ESS above 90%, and (ii) DR provides no advantage over Direct under low overlap, effectively reducing to the outcome model alone. These findings motivated the CLE bound and its TTC/AB diagnostics, which explain why high ESS is insufficient.
@@
    Does calibration work? Reward calibration should reduce RMSE, improve ranking accuracy, and with calibration-aware inference restore valid CI coverage.
@@ 5.1 Experimental design
Oracle $Y$: gpt-5-2025-08-07 (OpenAI, 2025) quality scores (0–100, normalized to $[0,1]$) collected for every response. Judge $S$: gpt-4.1-nano-2025-04-14 scores on all responses (temperature 0); approximately $16\times$ cheaper than oracle. [^2] Pricing as of November 2025 via OpenAI API. Note: The oracle model operates at temperature 1.0 (temperature 0 was not available for this model at experiment time), introducing stochasticity in oracle labels; we did not repeat-score responses, so coverage results may be conservative. All estimators except naive-direct calibrate judge scores using reward calibration. Base variants use monotone-only calibration (Z=SZ{=}S); +cov variants use two-stage calibration with response length as a covariate ($Z{=}g(S,\text{length})$). For DR methods, rewards $R=\hat{f}_{\mathrm{all}}(Z)$ use the full calibrator while the outcome model $\hat{q}=\hat{f}^{(-k)}(Z)$ uses cross-fitted (leave-one-fold-out) predictions; the difference $R-\hat{q}$ captures residual information for the IPS correction.
@@ 5.2 Results
Table 1: Accuracy & Uncertainty Metrics
Estimator $\mathrm{RMSE}^{\mathrm{d}}$ $\downarrow$ IS (interval score) $\downarrow$ Coverage
naive-direct 0.0828 2.9580 0.0 0.0039 90.9 79.6 0.817 0.4
direct 0.0223 0.0542 85.0 0.0068 91.8 84.1 0.837 0.7
direct+cov 0.0256 0.0568 86.1 0.0074 94.4 89.5 0.888 1.6
SNIPS 0.1596 0.7379 98.3 0.1815 38.3 8.7 -0.235 0.3
SNIPS+cov 0.1842 0.7544 97.8 0.1842 37.4 6.4 -0.252 0.3
calibrated-ips 0.0246 0.5727 99.1 0.0963 47.1 19.0 -0.059 0.3
calibrated-ips+cov 0.0262 0.6084 99.2 0.1027 46.5 20.5 -0.070 0.4
dr-cpo 0.0459 0.2203 99.2 0.0371 78.3 46.3 0.566 3.1
dr-cpo+cov 0.0570 0.2657 99.4 0.0449 79.1 50.1 0.581 9.6
calibrated-dr-cpo 0.0224 0.1357 99.5 0.0230 90.9 81.0 0.818 3.2
calibrated-dr-cpo+cov 0.0260 0.1398 99.3 0.0236 94.1 87.7 0.882 9.7
stacked-dr 0.0224 0.0788 95.6 0.0125 91.9 83.1 0.838 8.9
stacked-dr+cov 0.0281 0.0806 96.1 0.0128 92.1 84.8 0.842 38.2
$\downarrow$: lower is better, $\uparrow$: higher is better, $\to 95$: target is 95%. Bold: best (closest to target for Coverage %), underlined: second-best. Metrics averaged across all regimes.
@@
Yes, for both accuracy and inference. Table 1 shows calibration improves accuracy across method families. For Direct: uncalibrated naive-direct achieves 90.9% pairwise vs. calibrated direct+cov at 94.4%. For DR: raw-weight dr achieves 78.3% pairwise ($\tau{=}0.57$) vs. weight-stabilized calibrated-dr at 90.9% ($\tau{=}0.82$), a 12.6 percentage point gain (both use reward calibration; the gain is from weight stabilization). For inference: uncalibrated CIs (naive-direct) achieve 0% coverage: nominal 95% CIs capture the true value in 0/50 seeds, a failure mode independently replicated by Chen et al. (2026), who observe 0% coverage for naive estimation under judge bias. Bootstrap inference with bias-corrected estimation ($\hat{\theta}_{\mathrm{aug}}$) restores coverage: Direct achieves $\sim$}95% across oracle fractions (Table 3), as does stacked-DR. The calibration uncertainty share of total variance ranges from 0% (100% oracle) to over 50% (5% oracle), explaining why ignoring calibration uncertainty produces catastrophically narrow CIs (Figure 5).
Figure 4: CJE output: policy value estimates at n=1000n{=}1000, 25% oracle. Red diamonds show oracle ground truth; blue circles show CJE estimates with 95% CIs. CIs capture the true value for policies satisfying transportability; unhelpful (which violates transportability) shows slight miscoverage, a failure mode flagged by the transportability test (Figure 6).
@@
This was our main surprise. Despite weight stabilization boosting ESS from $1\%$ to $80\%$ for target policies, calibrated IPS ranking performance remains near-random: 47% pairwise accuracy vs. 50% chance baseline. We expected weight stabilization to be sufficient; it was not.
@@
We expected DR to outperform Direct by combining logged data with fresh draws. Instead, direct+cov slightly outperforms calibrated-dr+cov (94.4% vs. 94.1%). The explanation: under low overlap, DR’s IPS component contributes noise rather than information. DR estimators are consistent if either the propensity model or the outcome model is correct. Given IPS’s failure, DR effectively relies entirely on the outcome model $\hat{q}(Z)$, not the propensities. This is exactly the regime CLE identifies: TTC is low, so logged data contributes little. DR $\approx$ Direct ++ extra variance from unstable weights. Implication: When teacher forcing is unreliable, prefer Direct over DR; it achieves the same accuracy without requiring logprobs.
@@
Effect of Reward Calibration
@@
Direct + reward calibration 0.023 95% –72%
Effect of Weight Stabilization
@@
Calibrated DR, raw weights 0.046 99%
Calibrated DR + stabilization 0.023 100% –50%
Calibrated DR + stabilization + cov 0.026 99% –43%
@@
Direct + reward calibration + cov 0.026 95% –69%
Key findings: Reward calibration reduces RMSE by 72% and restores coverage (0%→95% with bootstrap). Weight stabilization reduces IPS RMSE by 84%. Both contribute additively.
Table 3: Coverage by UQ Method (target: 95%). Bootstrap with calibrator refit achieves near-nominal coverage; cluster-robust and jackknife-only methods undercover.
@@
Regime Cluster-Robust Cal-Aware Jackknife B=500B{=}500 B=2000B{=}2000
@@
Table 2 isolates the marginal contribution of each CJE component. Reward calibration reduces RMSE by 72% and is necessary to avoid catastrophic undercoverage: without it, 0% of CIs contain the true value; with bootstrap inference, coverage improves to $\sim$}95%. Weight stabilization reduces IPS RMSE by 84% while maintaining coverage, validating the $S$-monotone projection approach. Table 3 compares UQ methods for Direct mode: naive cluster-robust SEs achieve only 15–52% coverage; calibration-aware jackknife improves to 70–89%; bootstrap with $\hat{\theta}_{\mathrm{aug}}$ (AIPW-style residual augmentation) achieves $\sim$}95% by correcting both variance and bias from calibration uncertainty.
Table 4: Weight Diagnostics: Weight Stabilization. Raw SNIPS weights have catastrophic ESS ($1\%$) and heavy tails ($\alpha<2$); weight stabilization recovers ESS $80\%$ with light tails.
@@
Policy SNIPS Stab. SNIPS Stab. TTC CLE
@@
Figure 5: Uncertainty decomposition at n=1000n{=}1000 for direct+cov. Yellow: base sampling uncertainty (constant $\sim$}0.005). Orange: calibration uncertainty (dominates at low oracle fractions, contributing $\sim$}90% of total variance at 5% oracle; shrinks to $\sim$}30% at 100% because the estimator still fits a calibrator).
@@
Figure 6 applies the policy-wise mean transport test (4) to validate whether the base-trained calibration can be reused across target policies. We test $H_{0,\pi^{\prime}}:\mathbb{E}_{\pi^{\prime}}[Y-f(S,X)]=0$ for each policy with Bonferroni correction ($\alpha{=}0.0125$). Clone, Parallel Universe, and Premium all pass: their residual distributions are centered at zero, implying that Direct + calibrated surrogate remains mean-unbiased for these policies. Unhelpful fails (mean residual 0.31-0.31, $p<0.001$). By 4, this implies $V^{\mathrm{oracle}}_{\pi^{\prime}}-V^{\mathrm{sur}}_{\pi^{\prime}}\approx-0.31$: the surrogate systematically overestimates oracle quality for adversarial responses. This flags a failure of mean transportability: absolute value estimates for this policy are biased, triggering the REFUSE-LEVEL gate (App. D.6). Because Unhelpful’s bias (0.31-0.31) does not invert its rank relative to other policies (corrected value ${\approx}0.48$ remains well below ${\sim}0.75$), rankings remain valid; only absolute levels are refused. Practical implication: This test decides whether a single calibration can be safely reused across policies; if a policy fails, either recalibrate with policy-specific oracle data or fall back to oracle-only evaluation for that cell. Remark (transport correction). When transport fails, the estimated mean residual $\hat{\Delta}$ from the audit slice can serve as an intercept correction: $\hat{V}_{\mathrm{corr}}=\hat{V}_{\mathrm{sur}}+\hat{\Delta}$. However, we recommend refusing level claims by default, as the correction only fixes the mean and policy-specific recalibration is preferable when feasible.
Judge-policy confound.
When a model judges its own outputs, transport can fail due to self-preference bias. On MBPP code generation (Appendix G), GPT-4o-mini judging its own outputs showed a significant mean residual (0.065-0.065, p=0.006p{=}0.006). Never judge a model with itself without auditing transport; this failure mode cannot be fixed by calibration alone.
5.3 Comparison to Misclassification Correction
A natural alternative to continuous calibration is misclassification correction (Rogan & Gladen, 1978), recently adapted for LLM judges by Lee et al. (2025). Both approaches are calibration-based; they differ in what signal they use (continuous vs. thresholded) and in how uncertainty propagates. Lee et al. assume confusion-matrix stability—that $\Pr(\hat{Z}\mid Z)$ is constant across distributions—analogous to our transport condition.
Setup.
This comparison uses a separate experiment from the main benchmark, designed for an apples-to-apples evaluation with Lee et al.’s binary setting. We test on 9,999 Arena preference pairs (50.5% A-wins), the most favorable domain for binary methods: mean Youden’s $J=0.41$ across policies, compared to $J=0.05$ on code correctness where binary estimation is ill-conditioned. Both methods are evaluated on the same 3 target policies (verbose, terse, contrarian) with identical oracle budgets, test splits, seeds, and judge model (gpt-4o-mini). Confidence intervals follow Lee et al. (2025) Eq. 6 with Agresti-Coull adjustment.
Table 5: Misclassification correction vs. continuous calibration on Arena preferences (mean Youden’s $J\approx 0.41$). Both methods achieve near-nominal 95% CI coverage, but CJE produces $9\times$ narrower confidence intervals while achieving 93% lower RMSE even when Rogan-Gladen uses the optimal threshold.
Method RMSE 95% CI Coverage CI Half-Width
CJE (direct) $\mathbf{0.004}$ 96.4% $\mathbf{0.015}$
Rogan-Gladen ($\tau{=}0.5$) $0.063$ 95.6% 0.135
Rogan-Gladen ($\tau{=}0.6$, best) $0.061$ 98.9% 0.136
Setup: 9,999 Arena preference pairs (50.5% A-wins); 3 target policies (verbose, terse, contrarian); identical oracle budgets per cell. CIs follow Lee et al. (2025) Eq. 6. Threshold $\tau$ binarizes the continuous judge score; $\tau{=}0.5$ is the natural argmax, $\tau{=}0.6$ is optimal from a sweep over $\{0.3,0.4,0.5,0.6,0.7\}$.
Results.
Table 5 shows both methods achieve near-nominal coverage. However, even with the optimal threshold ($\tau{=}0.6$), CJE achieves 93% lower RMSE (0.004 vs. 0.061) and $9\times$ narrower confidence intervals (half-width 0.015 vs. 0.136).
Why the gap?
[…]
Implication.
Even when misclassification correction is numerically well-conditioned ($J$ well above zero), continuous calibration remains substantially more sample-efficient. Both approaches require transportability diagnostics: Lee et al.’s method relies on confusion-matrix stability; CJE relies on score transport. The key difference is that CJE makes diagnostics actionable—we gate the calibrated estimator on transport passing, preventing calibration from damaging estimates when the assumption fails. CJE also generalizes beyond preferences: on MBPP code generation with unit-test ground truth, calibration reduces RMSE by 32–84% for transport-valid policies at 10%+ oracle fraction (Appendix G).
Distribution shift robustness.
Lee et al. (2025) emphasize that their estimator is robust to shifts in $\Pr(Z)$ between calibration and test sets (their Section 7), since confusion-matrix parameters $\Pr(\hat{Z}\mid Z)$ do not depend on $\Pr(Z)$. We tested this claim by varying the calibration-set class balance while holding the test set at its natural rate ($\Pr(Z{=}1)\approx 0.51$). Results confirm their prediction: Rogan-Gladen bias remains near zero across all shifts (mean $|\mathrm{bias}|<0.007$), while CJE bias scales linearly with the calibration–test mismatch (0.14-0.14 at $\Pr(Z{=}1)_{\mathrm{calib}}{=}0.30$, +0.14+0.14 at $0.70$). This is a genuine advantage of confusion-matrix methods: when practitioners cannot control the calibration distribution, misclassification correction provides robustness that continuous calibration lacks. However, when calibration and test distributions match—the typical experimental design—CJE’s efficiency advantage dominates (RMSE $0.009$ vs. $0.032$ at no shift).
@@ Results.
Figure 7: Sample size planning for direct+cov. (A) RMSE by sample size and oracle fraction. (B) Standard errors including calibration uncertainty. (C) MDE contours showing the smallest effect detectable at 80% power. Dashed lines show common detection thresholds; cells with MDE <0.02<0.02 enable detecting 2-point quality differences. For a fixed label budget, spread labels across more samples (larger $n$, lower oracle %) rather than concentrating labels on fewer samples.
@@
Calibrated IPS Required + weight stabilization
@@ 5.4 Practical takeaways
    Default to Direct + two-stage calibration. For open-ended generation under nontrivial policy shifts, logs-only OPE faces two compounding issues: TF brittleness and coverage sparsity. Direct requires no overlap assumption and achieves the best ranking accuracy in our benchmark (Table 7).
@@
    Audit transport for high-stakes decisions. When reporting absolute values or making deployment decisions, run the mean transport test on a small oracle slice per target policy (or time window). Pass $\to$ reuse calibration; fail $\to$ recalibrate with policy-specific data or refuse level claims for that cell. For exploratory analysis or ranking-only comparisons, the audit can be deferred.
@@
    Gate on TTC before relying on logged data. If TTC <0.70<0.70, CLE implies a precision floor that weight stabilization cannot overcome; in our benchmark, only near-clone policies pass this threshold. For constrained settings (short outputs, near-clone shifts), OPE may remain viable; verify with diagnostics before trusting logged-data estimates.
@@
    Always use bootstrap inference with $\hat{\theta}_{\mathrm{aug}}$. Standard CIs severely under-cover (15–52%); calibration-aware jackknife improves but still falls short (70–89%); bootstrap with bias correction achieves near-nominal coverage ($\sim$}95%).
@@
    Use calibration uncertainty share to guide resource allocation. The fraction $\widehat{\mathrm{Var}}_{\mathrm{cal}}/\widehat{\mathrm{Var}}_{\mathrm{total}}$ identifies bottlenecks: if calibration uncertainty share $50\%$, collect more oracle labels; if calibration uncertainty share $20\%$, collect more evaluation prompts (Figure 5).
  6. 6.
@@
  7. 7.
    Optimize the budget ratio. How many oracle labels do you need? The optimal ratio of labels ($m$) to scores ($n$) follows a square-root law based on costs and variances (Appendix F). Check the calibration uncertainty share $\omega$ to equalize marginal utility: if $\frac{\omega}{1-\omega}>\frac{\mathrm{Spend}_{\mathrm{oracle}}}{\mathrm{Spend}_{\mathrm{surrogate}}}$, the experiment is under-investing in oracle labels.